OpenStudy (anonymous):

A bank randomly selected 240 checking account customers and found that 115 of them also had savings accounts at this same bank. Construct a 98% confidence interval for the true proportion of checking account customers who also have savings accounts. (Give your answers correct to three decimal places.)

OpenStudy (anonymous):

phat +/- E E= z (alpha/2) * sqrt ( (phat*qhat)/n ) z (alpha/2) = 2.33 phat = 115/240 qhat = 125/240 n=240 E=.0751 phat - E = .404 phat + E = .554 .404<p<.554